If Ms. Sarmiento is investing Php 200,000.00 with a simple interest rate of 3.5% for 8 years, how much will the money of Ms. Sarmiento be after 8 years?
Answer by ikleyn(52788)

.

Use the formula for the future value of the simple interest account

    future value = 200000*(1+0.035*8) = 200000*1.28 = 256000 Php.    ANSWER

Solved.

------------------

To see many other similar  (and different)  solved problems on simple interest accounts,  look into the lesson
    - Simple interest percentage problems
in this site.

Learn the subject from there.